OpenStudy (anonymous):

$6,000 dollars is invested in two different accounts earning 3% and 5% interest. At the end of one year, the two accounts earned $220 in interest. How much money was invested at 3%? $2,000 $3,000 $4,000

hero (hero):

Hint: x + y = 6000 .03x + .05y = 220

OpenStudy (anonymous):

thanks... your my hero

OpenStudy (anonymous):

lol xD

hero (hero):

If you can let me know what the answer is, I'll give you a medal.

OpenStudy (anonymous):

okay(: , i'm trying to set it up... .03x+.05y=220 then i moved y to the other side like .03x=220-.05y then i got x=7333.3333-1.666y

OpenStudy (anonymous):

but i have a feeling i did something wrong

OpenStudy (anonymous):

am i doing it right?

hero (hero):

Not exactly

OpenStudy (anonymous):

what am i doing wrong?

hero (hero):

Multiply the second equation by 100

hero (hero):

Let me know what you get afterwards

OpenStudy (anonymous):

.03x=220-.05y is this what you mjean by second equation?

hero (hero):

1st Equation x + y = 6000 2nd Equation .03x + .05y = 220 I don't understand what you're confused about here.

hero (hero):

Multiply the second equation by 100 and let me know what you get afterwards

OpenStudy (anonymous):

3x-5y=22000

OpenStudy (anonymous):

3x-y=4400

OpenStudy (anonymous):

good so far?